OpenStudy (anonymous):

(-1, 3) and perpendicular to y = -one thirdx + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (mathmale):

Taylor: It'd be helpful if you'd please write out all of the directions. I believe you mean the following: Find the equation of the straight line that passes through (-1,3) which is also perpendicular to the line \[y=\frac{ -1 }{ 3 }x +7.\] All right. If the new line is to be perpendicular to the above line, what is the slope of the new line? In other words, given a line with a known slope, how do we find the slope of a line perpendicular to the given line? Once you have the slope of the new line, I'd suggest you use the point-slope form of an equation for a straight line: \[y-y _{1}=m(x-x _{1}).\] Good luck!

OpenStudy (anonymous):

If two lines are perpendicular and have slopes m1 and m2 ,then m1*m2=-1

OpenStudy (texaschic101):

do you know the slope Tayloresc ?

OpenStudy (texaschic101):

In y = mx + b, the slope is in the m position. So, in y = -1/3x + 7, what is the slope ?

OpenStudy (texaschic101):

let me know when you find the slope